### v1

# Decoupled Generative Modeling for Human-Object Interaction Synthesis

Abstract
Synthesizing realistic human-object interaction (HOI) is essential for 3D computer vision and robotics, underpinning animation and embodied control. Existing approaches often require manually specified intermediate waypoints and place all optimization objectives on a single network, which increases complexity, reduces flexibility, and leads to errors such as unsynchronized human and object motion or penetration. To address these issues, we propose Decoupled Generative Modeling for Human-Object Interaction Synthesis (DecHOI), which separates path planning and action synthesis. A trajectory generator first produces human and object trajectories without prescribed waypoints, and an action generator conditions on these paths to synthesize detailed motions. To further improve contact realism, we employ adversarial training with a discriminator that focuses on the dynamics of distal joints. The framework also models a moving counterpart and supports responsive, long‑sequence planning in dynamic scenes, while preserving plan consistency. Across two benchmarks, FullBodyManipulation and 3D-FUTURE, DecHOI surpasses prior methods on most quantitative metrics and qualitative evaluations, and perceptual studies likewise prefer our results.
**footnotetext: Corresponding author.
1 Introduction
Realistic human-object interaction synthesis (HOI) is a fundamental task with broad impact in computer vision and robotics [bhatnagar2022behave, xu2025interact, liu2022hoi4d, cao2023detecting, chen2023detecting]. These capabilities form the basis of modern 3D systems, enabling human motion animation and humanoid control [zhu2023human, xue2025human, liu2025core4d]. Synthesizing interaction requires perceiving the pose of the object and awareness of the target goal, followed by the generation of a plausible sequence of human joint configurations that performs the required manipulation safely and intentionally [antoun2023human, yang2024lemon, yang2024f, xu2025intermimic]. Recent approaches [li2024controllable, wu2025human] rely on interpretable natural language instructions to specify tasks, yet producing interactions that are faithful to the prompt while keeping realism and diversity remains a difficult challenge.
Figure 1: Overview of DecHOI for dynamic human-object interaction synthesis. The framework decouples trajectory planning and interaction synthesis, enabling collision detection and responsive re-planning for realistic, contact-consistent motion.
Prior methods [li2024controllable, wu2025human] generate 4D motion sequences of humans and objects by conditioning a diffusion model on path waypoints and text instructions. While these methods provide a strong foundation for interaction modeling that is capable of producing desired behaviors, this design introduces two challenges: i) reliance on specified intermediate waypoints that the human must follow, and ii) high optimization complexity from assigning the entire objective to a single denoising network. Specifically, for the first challenge, at inference time, the user must manually provide not only the start and goal points but also intermediate waypoints. Such reliance on externally supplied constraints narrows the model’s generative scope and reduces autonomy, while introducing procedural burden for the user. For the second challenge, HOI requires the model to simultaneously resolve the motion of both the human and the object in every frame. Assigning all objectives to a single network elevates optimization complexity and risks practical intractability. The resulting complexity often produces unsynchronized interactions, such as hovering or penetration of objects, which harms plausibility and degrades overall quality. These observations motivate us to design a new intermediate representation that is simple, flexible, informative, and still enables interactions that remain faithful to the instruction and plausible for HOI generation.
To this end, we propose Decoupled Generative Modeling for Human-Object Interaction Synthesis (DecHOI), a novel framework that generates trajectories and motions as separate processes. In DecHOI, a trajectory generator (TG) first produces human and object trajectories, and an action generator (AG) then synthesizes detailed actions conditioned on the generated paths. This decoupling allows the TG to produce diverse and accurate paths without reliance on specified waypoints, while the AG focuses exclusively on human actions and object motion, capturing more fine-grained details. Partitioning the model into two lightweight expert networks reduces the optimization complexity for each branch and mitigates unsynchronized interactions.
Another key factor for human-level interaction is precise hand and foot control [hao2024hand, mao2022contact]. To improve the robustness of hand and foot contact and reduce undesired penetration, we introduce adversarial training applied to the object and to the distal joints. A compact discriminator focuses on hand-object interaction signals and foot dynamics, shaping the learning distribution and enabling realistic control.
For applications, the previous model [li2024controllable] enables 3D scene-aware long-sequence interaction synthesis but remains restricted to static environments. To overcome this limitation, our framework models a moving counterpart and equips the human agent to either avoid the counterpart or wait, allowing it to respond to dynamic environments. This improves practicality in dynamic scenes by aligning actions with the plan and adapting to moving obstacles.
To validate the effectiveness of DecHOI, we evaluated our method on the FullBodyManipulation [li2023object] and observed superior performance on most quantitative metrics and qualitative assessments. We also report results on unseen objects from the 3D-FUTURE [fu20213d], demonstrating robustness and generalizability. For long-sequence interaction synthesis, we present scenarios in indoor environments furnished with diverse objects, where the human agent interacts with a moving counterpart and reaches various goal points by either avoiding the counterpart or waiting. The summary of our contributions is as follows.
  • We propose DecHOI, a decoupled framework that separates trajectory generation from fine‑grained action synthesis, reducing optimization complexity and removing the need for manual intermediate waypoints.
  • We improve coordination between distal joints and an object through adversarial training with a compact discriminator, thereby reducing interpenetration.
  • We enable responsive planning with a long-sequence planner that dynamically adapts to moving counterparts, supporting scene-aware interaction.
  • Experiments conducted on various benchmarks demonstrate that DecHOI achieves state-of-the-art performance, surpassing previous methods in terms of realism, accuracy, and diversity.
2 Related Work
2.1 Human Action Generation
Human action generation aims to synthesize realistic 3D motion sequences from conditioning signals such as text descriptions, action labels, and joint-level constraints [meng2025rethinking, zhang2025energymogen, fan2024freemotion, zou2024parco, jiang2023motiongpt]. T2M [guo2022generating] samples a motion length from text and then synthesizes the motion with a temporal VAE over motion snippet codes. MDM [tevet2023human] adapts classifier-free diffusion [ho2022classifier] with a transformer backbone for high-fidelity motion, and PriorMDM [shafir2024human] treats a pre-trained diffusion model as a generative prior to enable controllable composition. ACMDM [meng2025absolute] uses absolute global joint coordinates in a streamlined transformer diffusion model, improving fidelity and text alignment. OmniControl [xie2024omnicontrol] adds spatial constraints for any joint at any time using analytic guidance with a learned realism prior, while MoMask [guo2024momask] enables text-guided temporal inpainting via masked generation with hierarchical motion tokens. However, these works primarily address single human motion and neither manipulate objects nor perceive or adapt to the surrounding scene [wang2025hsi, wang2024move, li2024genzi]. In contrast, we target scene-aware human-object interaction that avoids obstacles and manipulates objects in accordance with the specified intent.
2.2 Human-Object Interaction Synthesis
Human-object interaction synthesis (HOI) takes text instructions or object state as input and generates purposeful 3D human-object motion [xue2025guiding, jia2025primhoi, zhang2025interactanything, song2024hoianimator]. For example, OMOMO [li2023object] conditions on full per-frame object motion and enforces contact by integrating hand kinematics during whole body generation. CHOIS [li2024controllable] and HOIFHLI [wu2025human] show that per-frame object conditioning is unnecessary because language instruction paired with sparse 3D object waypoints, together with the initial human and object states, stabilizes intended behaviors. As a complementary effort, CG-HOI [diller2024cg] jointly generates human and object motion with contact cues, using contact-guided conditioning to improve realism. Despite these gains, a single network that models trajectory, pose, and contact is still hard to optimize and often produces desynchronization and contact artifacts. We instead decouple trajectory and action generation and apply contact-aware adversarial training, yielding robust synchronized HOI.
Figure 2: Architecture of DecHOI showing the decoupled trajectory and action generation process. Conditioned on the text instruction, geometry, current human and object poses, and a goal point, the trajectory generator plans paths, while the action generator produces joint motions on these paths to yield synchronized, contact-aware interactions. The right panels detail the Trajectory and Action Generators.
2.3 Path-conditioned HOI
Path-conditioned HOI first specifies a long-term trajectory via waypoints, object paths, or scene-level guidance, and then synthesizes it at the motion level [yi2024generating, cen2024generating, lee2023locomotion]. CHOIS generates synchronized long-horizon human-object motion in 3D scenes by conditioning diffusion on sparse object waypoints obtained from Habitat [savva2019habitat, szot2021habitat] for the given start and target positions. NIFTY [kulkarni2024nifty] couples an object-conditioned motion diffusion model with a learned interaction field that evaluates distances to a feasible interaction manifold and guides sampling toward contact-consistent behaviors. OMOMO [li2023object] takes per-frame object motion data as input and synthesizes manipulation by denoising, incorporating hand motion to maintain contact during whole body generation. However, in inference, these formulations retain a fixed plan or guidance and therefore do not adapt to external scene changes. By contrast, our method updates the plan, enabling adaptation to external scene changes while preserving high-level intent.
3 Method
Given an instruction, current human and object poses, and a goal point, our framework synthesizes an interaction that moves the object to the goal as instructed. As illustrated in Fig. 2, we adopt a decoupled generation architecture with a trajectory generator and an action generator (Sec. 3.2) to reduce reliance on explicit waypoints and simplify the optimization objective. To improve contact realism and reduce interpenetration, we employ adversarial training with a compact discriminator that focuses on the dynamics of distal joints (Sec. 3.3). In Sec. 3.4, we present the overall optimization procedure and the inference-time guidance strategies for our method.
3.1 Background: Denoising Diffusion Model
Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising, tevet2023human] are generative models that learn to approximate complex data distributions by gradually adding noise and then removing it. This paradigm has become dominant in recent work on human action generation [tevet2023human, shafir2024human, meng2025absolute, xie2024omnicontrol] and interaction synthesis [li2023object, li2024controllable, wu2025human, xu2023interdiff]. The forward process progressively corrupts a clean sample $\mathbf{x}_{0}$ by adding Gaussian noise over $N$ steps, defined as:
$q(\mathbf{x}_{n}\mid\mathbf{x}_{0})=\mathcal{N}\!\bigl(\mathbf{x}_{n};\sqrt{\bar{\alpha}_{n}}\mathbf{x}_{0},(1-\bar{\alpha}_{n})\mathbf{I}\bigr),$ (1)
where $\alpha_{n}=1-\beta_{n}$ and $\bar{\alpha}_{n}=\prod_{s=1}^{n}\alpha_{s}$. The forward process is a Markov chain with Gaussian transitions. As $n$ increases, $\mathbf{x}_{n}$ converges to an isotropic Gaussian. The reverse process learns to invert this noising procedure with a parameterized model $p_{\theta}$ that progressively denoises $\mathbf{x}_{n}$ to recover clean data:
$p_{\theta}(\mathbf{x}_{n-1}\mid\mathbf{x}_{n},\mathbf{c})=\mathcal{N}\!\bigl(\mathbf{x}_{n-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{n},n,\mathbf{c}),\sigma_{n}^{2}\mathbf{I}\bigr),$ (2)
where $\mathbf{c}$ denotes conditioning inputs such as text embeddings, object geometry, or start and goal states. Following prior formulations [li2024controllable, wu2025human], our denoising network predicts the clean sequence rather than the noise:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},n\sim[1,N]}\|\hat{\mathbf{x}}_{0}-\mathbf{x}_{0}\|_{1}.$ (3)
Predicting $\hat{\mathbf{x}}_{0}$ has been found to yield better results for motion data and allows guidance losses to be applied at each denoising step.
3.2 Decoupled Generative Modeling
Jointly estimating and optimizing human and object poses is non-trivial for a single network to handle. We address this challenge with decoupled generative modeling that separates trajectory generation from action generation. At the start of the pipeline the inputs are the object pose sequence $P_{o}\in\mathbb{R}^{T\times 12}$ with $T$ frames, where each frame contains the global 3D position and a 9-parameter relative rotation matrix, and the human pose sequence $P_{h}\in\mathbb{R}^{T\times D_{h}}$, where $D_{h}$ comprises global joint coordinates and per-joint 6D rotation parameters. To ensure accurate interaction and instruction following, inputs also include object geometry $B\in\mathbb{R}^{1024\times 3}$ represented as a Basis Point Set (BPS) [prokudin2019efficient] and a text instruction.
Trajectory Generator. We formulate the trajectory generator (TG) as a conditional denoising diffusion model [ho2020denoising]. Given an input sequence, we apply a forward noising process for $N$ steps following a Markov chain. For conditioning, the human and object poses of the start frame remain clean, and the object position in the end frame is also kept clean to set the goal point. Then these noisy data are concatenated with the geometry embedding $F_{\text{obj}}\in\mathbb{R}^{D}$, computed by a simple MLP encoder from $B$. We obtain the conditioning input $X\in\mathbb{R}^{T\times(12+D_{h}+D)}$ for the denoising network. A transformer-based denoising network of TG generates both global object and human trajectories that are faithful to the instruction. For example, given the instruction “Lift the chair, move the chair, and put down the chair,” the object trajectory rises in the early part of the sequence and descends as it approaches the goal point. To learn this alignment, we incorporate a text embedding $F_{\text{text}}\in\mathbb{R}^{D}$ obtained from a CLIP encoder [radford2021learning] into the model inputs. In contrast to prior methods [li2024controllable, wu2025human] that condition by concatenating the embedding along the sequence dimension, we use a cross-attention layer to apply the conditioning explicitly, which yields stronger alignment between text and motion features. Afterward, the TG yields continuous 3D paths for the object $\hat{\mathcal{T}}_{o}\in\mathbb{R}^{T\times 3}$ and the human $\hat{\mathcal{T}}_{h}\in\mathbb{R}^{T\times 3}$.
Action Generator. Once reliable human and object trajectories have been generated, we use them as richer conditions. The action generator (AG) receives the same input $X$ as the trajectory generator (TG), except that the noisy global positions of the object and the human root in all $T$ frames are replaced with the trajectories produced by TG. This dense conditioning provides stronger priors and reduces the complexity of the learning objective. AG is a lightweight diffusion model that employs a transformer-based denoising network with the same architecture as TG. Given the conditioned noisy input, it generates the full pose of the object $\hat{P}_{o}\in\mathbb{R}^{T\times 12}$ and the pose of the human joints $\hat{P}_{h}\in\mathbb{R}^{T\times D_{h}}$. The human pose generated is then used for parametric human modeling, and we apply SMPL-X [pavlakos2019expressive] to reconstruct the human mesh and pose.
3.3 Adversarial Training for Distal Joints
Realistic human-object interaction (HOI) requires not only plausible motion but also reliable contact [cseke2025pico, gu2024contactgen, jiang2024scaling]. Language instructions such as “Lift the chair” or “Kick the box” rely on distal joints in the hands and feet. Existing models often cause the hands and feet to drift into empty space or intersect objects [yuan2023physdiff, xu2025interact, mao2022contact]. To address these issues, we introduce an adversarial training mechanism that regularizes contact by focusing on the distal joints.
In general, the most reliable cue to decide whether an input is real or fake is the distance between the object surface and the distal joints [mao2022contact, cseke2025pico, huang2022capturing, tripathi2023deco]. In ground truth, these distances are small due to complete contact, whereas generated results often exhibit larger gaps. Optimizing the generator so that the discriminator cannot distinguish real from fake therefore acts as a regularizer that drives contact toward completeness. Motivated by this observation, we design a compact discriminator $\mathcal{D}$ that differentiates real and fake using distal joint kinematics together with object geometry. As shown in Fig. 3, $\mathcal{D}$ receives global coordinates of the hands $H\in\mathbb{R}^{T\times 2\times 3}$ and feet $F\in\mathbb{R}^{T\times 2\times 3}$ over all $T$ frames, obtained by forward kinematics from the 6D joint rotations. In addition, we sample $M$ points from the geometry of the object $B$ to obtain $B’\in\mathbb{R}^{T\times M\times 3}$ and apply the relative rotation and translation for each frame. For real data, we feed clean inputs $x$, and for fake data, we use the outputs $\hat{x}$ produced by AG. The discriminator is trained with the following loss:
$\begin{split}\mathcal{L}_{\mathcal{D}}&=\frac{1}{T}\sum_{t=1}^{T}\big([1-s^{(r)}_{t}]_{+}+[1+s^{(f)}_{t}]_{+}\big),\\ s^{(r)}_{t}&=\mathcal{D}(x)_{t},\quad s^{(f)}_{t}=\mathcal{D}(\hat{x})_{t},\end{split}$ (4)
where $s^{(r)}_{t}$ and $s^{(f)}_{t}$ denote the scores for real and fake data at frame $t$. After completing the discriminator training stage, we train AG to fool the discriminator by minimizing the following generator objective:
$\mathcal{L}_{G}=-\frac{1}{T}\sum_{t=1}^{T}s^{(f)}_{t}$. (5)
Through adversarial training, the network enforces higher contact fidelity and synthesizes realistic interactions.
Figure 3: Adversarial module of DecHOI, where a hand and foot-focused discriminator contrasts real and generated interactions to enhance contact realism.
Condition Matching Human Motion Quality Interaction Quality GT Difference
[…]
Lin-OMOMO [li2023object] - - 10.04 0.44 0.45 16.81 6.65 0.75 0.56 0.59 0.70 0.65 17.45 29.20 75.65 -
Pred-OMOMO [li2023object] 2.34 9.66 6.99 0.38 0.63 13.01 6.68 0.67 0.48 0.52 0.59 0.60 23.63 49.12 51.86 1.23
CHOIS [li2024controllable] 1.92 8.01 6.15 0.41 0.68 1.58 8.31 0.74 0.53 0.58 0.66 0.61 18.86 44.04 51.86 1.23
HOIFHLI [wu2025human] 1.73 7.65 4.77 0.38 0.62 2.06 8.55 0.77 0.61 0.64 0.58 0.61 19.31 40.87 50.96 1.18
Ours (DecHOI) 1.59 6.91 4.42 0.38 0.72 0.33 8.86 0.80 0.64 0.67 0.53 0.54 15.27 25.47 22.96 0.86
Table 1: Quantitative comparison on the FullBodyManipulation [li2023object] with CHOIS [li2024controllable], HOIFHLI [wu2025human], and OMOMO [li2023object] variants (Lin-OMOMO and Pred-OMOMO) across four categories of evaluation metrics. Arrows indicate direction: ($\uparrow$) means higher is better, ($\downarrow$) means lower is better, and ($\rightarrow$) means closer to the real data value is better. The real-data DIVDIV reference is 9.02.
Condition Matching Human Motion Quality Interaction Quality
Methods $T_{s}\downarrow$ $T_{e}\downarrow$ $H_{\text{feet}}\downarrow$ $FS\downarrow$ $R_{prec}\uparrow$ $FID\downarrow$ $C_{\%}$ $P_{\text{hand}}\downarrow$
Lin-OMOMO [li2023object] - - 7.92 0.48 0.57 8.85 0.26 0.20
Pred-OMOMO [li2023object] 4.72 10.92 6.61 0.45 0.59 5.16 0.40 0.17
CHOIS [li2024controllable] 5.75 10.28 4.20 0.42 0.61 2.04 0.46 0.18
Ours (DecHOI) 4.27 8.43 4.06 0.41 0.69 1.01 0.48 0.15
Table 2: Quantitative results on the 3D-FUTURE [fu20213d]. DecHOI achieves better trajectory accuracy, motion stability, and contact realism than CHOIS [li2024controllable] and OMOMO [li2023object] baselines.
3.4 Training and Inference
Objectives for Training. During optimization, the trajectory generator (TG) and the action generator (AG) are trained with distinct objectives. TG focuses exclusively on planning trajectories. Given the input $X$, it produces $\hat{\mathcal{T}}_{0}=\{\hat{\mathcal{T}}_{o},\hat{\mathcal{T}}_{h}\}$ and is trained to reconstruct the clean trajectory representation $\mathcal{T}_{0}=\{\mathcal{T}_{o},\mathcal{T}_{h}\}$ with an $L_{1}$ objective. The loss is defined as:
$\mathcal{L}_{\text{TG}}=\mathbb{E}_{\mathcal{T}_{0},n\sim[1,N]}\|\hat{\mathcal{T}}_{0}-\mathcal{T}_{0}\|_{1}.$ (6)
AG receives the same input $X$ as TG, while the trajectories are kept clean. AG is trained to reconstruct the entire motion $P_{0}=\{P_{o},P_{h}\}$ from which the generator predicts $\hat{P}_{0}=\{\hat{P}_{o},\hat{P}_{h}\}$. The reconstruction objective is:
$\mathcal{L}_{\text{AG}}=\mathbb{E}_{P_{0},n\sim[1,N]}\|\hat{P}_{0}-P_{0}\|_{1}.$ (7)
To further stabilize the reconstruction of the distal joints, we add a forward kinematic loss. Using the predicted relative joint rotations, we compute global hand positions $\hat{H}_{0}$ and global foot positions $\hat{F}_{0}$ from the pelvis root. These are supervised by clean distal joint positions $H_{0}$ and $F_{0}$ with:
$\mathcal{L}_{\text{FK}}=\|\hat{H}_{0}-H_{0}\|_{1}+\|\hat{F}_{0}-F_{0}\|_{1}.$ (8)
The total objective for training the generators combines the reconstruction and adversarial terms:
$\mathcal{L}=\lambda_{\text{TG}}\mathcal{L}_{\text{TG}}+\lambda_{\text{AG}}\mathcal{L}_{\text{AG}}+\lambda_{\text{FK}}\mathcal{L}_{\text{FK}}+\lambda_{\text{G}}\mathcal{L}_{\text{G}},$ (9)
where $\lambda_{\mathrm{TG}}$, $\lambda_{\mathrm{AG}}$, $\lambda_{\mathrm{FK}}$, and $\lambda_{\mathrm{G}}$ are scalar weights that balance the contributions of the terms.
Inference-time Guidance. Inspired by previous work [li2024controllable], our model applies reconstruction guidance to regularize the generation process. This design injects constraints at each denoising step without retraining the network for a specific purpose, which permits flexible control over the outputs. The process is formally represented as:
$\tilde{P}_{0}=\hat{P}_{0}-\alpha\Sigma_{n}\nabla_{P_{n}}\mathcal{F}(\hat{P}_{0}),$ (10)
where $\mathcal{F}$ is a regularization objective and $\alpha$ controls the perturbation strength. In our implementation, $\mathcal{F}$ encourages precise contact by penalizing distances between distal joints and the object surface, and stabilizes stance by minimizing deviations of foot joints from the ground plane to prevent hovering and penetration. Additional details are provided in the supplementary material.
4 Experiments
4.1 Datasets
Our experiments utilize two datasets designed to capture realistic human-object interactions across diverse indoor environments, following the evaluation setup in prior work [li2024controllable]. i) FullBodyManipulation [li2023object]: Comprising about 10 hours of synchronized human-object motion sequences, this dataset covers 15 distinct rigid objects. We consider only rigid body interactions and therefore exclude sequences with articulated objects. Motion data from 15 subjects are used for training, with data from two additional subjects reserved for testing, ensuring a consistent evaluation protocol. ii) 3D-FUTURE [fu20213d]: Containing a broad range of 3D furniture models. This dataset offers substantial geometric diversity. To evaluate generalization to unseen appearances and shapes, we substitute the 17 test objects in FullBodyManipulation with unseen 3D-FUTURE models of the same categories, enabling systematic assessment on previously unobserved instances.
4.2 Evaluation Metrics
Condition Matching.
This metric evaluates how accurately the generated motion aligns with the specified object conditions. We measure the Euclidean distance between the predicted and target object start and end positions $T_{s}$ and $T_{e}$ in the scene, reported in centimeters (cm).
Human Motion Quality.
We evaluate realism and plausibility with five metrics: foot height ($H_{\text{feet}}$), foot sliding ($FS$), R-precision ($R_{\text{prec}}$), Fréchet Inception Distance (FIDFID[heusel2017gans], and Diversity ($DIV$). $H_{\text{feet}}$ is the mean foot-to-ground height. $FS$ measures horizontal foot displacement during stance, following [li2024controllable]. Both are reported in centimeters. $R_{\text{prec}}$ (top-3) scores text-motion alignment. $FID$ compares generated and ground-truth motion distributions in a learned feature space. $DIV$ quantifies variation across samples under the same condition.
Interaction Quality.
Figure 4: Qualitative comparison of DecHOI with CHOIS [li2024controllable] and HOIFHLI [wu2025human] on the FullBodyManipulation [li2023object]. DecHOI produces stable contacts, smooth motion, and accurate object trajectories, while prior methods show drift, penetration, or inconsistent coordination between human and object motions.
This category evaluates interaction accuracy and naturalness. We report hand-object contact precision ($C_{\text{prec}}$), recall ($C_{\text{rec}}$), F1 ($C_{F1}$), and percent ($C_{\%}$) as an overall contact reliability measure. We also report hand-object ($P_{\text{hand}}$) and body-object ($P_{\text{body}}$) penetration to quantify interpenetration depth ($cm$).
Ground Truth Difference.
This metric quantifies the deviation of generated motions from ground truth motions in both spatial and rotational domains. We report mean per-joint position error (MPJPE), root translation error ($T_{\text{root}}$), and object translation error ($T_{\text{obj}}$) as Euclidean distances in centimeters. Orientation consistency is evaluated with object orientation errors ($O_{\text{obj}}$), computed as the Frobenius norm of rotation differences.
4.3 Quantitative Analysis
We quantitatively evaluate our approach against CHOIS [li2024controllable], HOIFHLI [wu2025human], and two OMOMO [li2023object] variants (Lin- and Pred-OMOMO) implemented following CHOIS on two datasets: FullBodyManipulation [li2023object] and 3D-FUTURE [fu20213d]. Because our method does not require specified intermediate waypoints, we evaluate all baselines without intermediate waypoint inputs to ensure a fair comparison. Note that Lin-OMOMO uses ground truth for all object-related signals, so metrics that are not applicable under this setting are reported as (–).
FullBodyManipulation.
As shown in Tab. 1, our model surpasses all baselines on most metrics. OMOMO [li2023object] variants condition on object pose at every frame, yet under our extremely limited inputs they underperform in overall human motion and interaction quality. Methods relying on sparse object conditioning [li2024controllable, wu2025human] also degrade on contact-related measures once waypoints are removed, as the reduced conditioning increases task complexity and leads to unsynchronized interactions. In contrast, DecHOI simplifies the problem through decoupled modeling, allowing the trajectory generator to produce accurate paths with low condition matching error and to attain competitive DIV with reduced reliance on waypoints. The resulting interactions exhibit higher realism and stronger alignment with instructions, as reflected by improved FID and $R_{prec}$. Although contact accuracy and penetration scores often trade off against each other, our adversarial training achieves balanced performance across both and produces realistic interactions. Further analyses and ablations are provided in the supplementary material.
3D-FUTURE.
To further evaluate the generalization capability of our approach, we test it on unseen 3D-FUTURE samples that share the same object categories as those in the FullBodyManipulation test set. As shown in Tab. 2, our method maintains consistent trends across metrics, achieving strong trajectory accuracy and high contact reliability. Moreover, the improved $FID$ indicates better alignment between generated and GT motion distributions. These results demonstrate that DecHOI effectively synthesizes realistic and coherent human-object interactions even for previously unseen objects, highlighting its robust generalization and broad applicability across diverse objects.
4.4 Qualitative Analysis
In Fig. 4, we present qualitative comparisons among CHOIS [li2024controllable], HOIFHLI [wu2025human], and our DecHOI on the FullBodyManipulation [li2023object]. As illustrated in the four scenes, DecHOI consistently produces more stable and realistic interactions. For the largebox scene, CHOIS and HOIFHLI yield unstable object motion, leading to box hovering and hand-object penetration. In contrast, DecHOI maintains a firmly grounded trajectory and cleaner contacts through stable global path modeling. In the plasticbox scene, baselines exhibit severe penetration between the human body and the object, along with misaligned hand contacts. In contrast, our adversarial training on distal joints mitigates these failures and enhances contact realism. For the tripod scene, both struggle with fine manipulation as the object drifts and rotates inconsistently, whereas our decoupled framework suppresses such unsynchronized artifacts and achieves more coherent motion. Lastly, in the woodchair scene, prior works show noticeable mesh intersection and incomplete contact, while DecHOI produces stable contacts and precise spatial alignment between the hand and object.
Figure 5: Qualitative comparison of DecHOI and CHOIS [li2024controllable] on the 3D-FUTURE [fu20213d], showing generalization to unseen objects.
As shown in Fig. 5, we further compare CHOIS and DecHOI on two unseen objects from the 3D-FUTURE  [fu20213d]: largetable and woodchair. CHOIS often fails to adapt to these novel geometries, leading to human-object intersection and motion instability, whereas our method preserves accurate contacts and smooth motion, demonstrating strong generalization to unseen shapes. More qualitative results and analysis are provided in the supplementary material.
Additionally, Fig. 6 visualizes the loss landscape to compare the training objective complexity of CHOIS, which uses a single network, and DecHOI. The surface of CHOIS appears noisy and highly rugged with many local minima, indicating sensitivity to initialization and unstable training. In contrast, DecHOI exhibits a smoother landscape with fewer local minima and more stable convergence. These observations suggest that our decoupled modeling effectively reduces overall optimization complexity.
4.5 Long-term Dynamic Planning
To evaluate long-term and dynamic human-object interactions, we introduce DynaPlan, which supports responsive planning and scene-aware interaction generation under multi-agent conditions. We equip both the agent and the moving counterpart with circular influence radii and detect potential collisions when their regions intersect. When a collision is detected, we re-plan with A*, adaptively choosing either a short detour or waiting to maintain goal-directed and collision-free motion while accounting for the future path of the counterpart. To handle uncertainty in that future path, we predict it with a pre-trained trajectory prediction network [mohamed2020social]. We evaluate performance on about 190 long-sequence indoor scenarios. Further details are provided in the supplementary material.
Quantitative Results.
Figure 6: Visualization of training loss landscapes for DecHOI and CHOIS [li2024controllable], demonstrating reduced optimization complexity.
Condition Matching Human Motion Quality Interaction Quality
Methods $T_{s}\downarrow$ $T_{e}\downarrow$ $H_{\text{feet}}\downarrow$ $FS\downarrow$ $C_{\%}$ $P_{\text{hand}}\downarrow$ $P_{\text{body}}\downarrow$ $P_{\text{o$\rightarrow$}s}\downarrow$ $P_{\text{h$\rightarrow$}s}\downarrow$
CHOIS [li2024controllable] 2.19 8.05 3.14 0.43 0.78 0.73 0.73 1.00 0.83
Ours (DecHOI) 1.90 7.98 2.85 0.45 0.76 0.65 0.63 0.72 0.65
Table 3: Quantitative results between CHOIS [li2024controllable] and DecHOI for responsive long-term interaction synthesis on the DynaPlan.
Tab. 3 summarizes the quantitative results for long-sequence human-object interaction synthesis on the DynaPlan. Compared to CHOIS [li2024controllable], DecHOI achieves lower trajectory errors $T_{s}$ and $T_{e}$, along with reduced instability, resulting in smoother and more stable motion over extended sequences. For interaction quality, DecHOI shows clear improvements across all penetration metrics, including $P_{\text{hand}}$, $P_{\text{body}}$, $P_{\text{o}\rightarrow\text{s}}$, and $P_{\text{h}\rightarrow\text{s}}$. Here, $P_{\text{hand}}$ and $P_{\text{body}}$ measure the penetration depth between the human mesh and manipulated objects, while $P_{\text{o}\rightarrow\text{s}}$ and $P_{\text{h}\rightarrow\text{s}}$ quantify collisions between the object or human and the surrounding static scene. Lower values on most metrics indicate that DecHOI effectively minimizes interpenetration and maintains collision-free trajectories even in cluttered indoor environments.
Figure 7: Visualization of DecHOI in long-sequence dynamic environments. The human agent (blue) adaptively re-plans its path when encountering a moving obstacle (green), choosing between detour and waiting behaviors to maintain goal-directed, collision-free motion.
These performance gains stem from the decoupled structure of the trajectory and action generators, which separates global path planning from detailed motion synthesis, combined with contact-aware adversarial learning that enforces spatial consistency throughout long-horizon interactions. Overall, DecHOI demonstrates superior stability, plausibility, and contact accuracy in dynamic multi-agent settings, outperforming CHOIS quantitatively.
Qualitative Results.
Fig. 7 visualizes DecHOI operating in two long-term dynamic indoor scenes from DynaPlan. In both scenes, the agent encounters a moving counterpart along its original path and initiates reactive re-planning to maintain safe, goal-directed motion. Throughout these sequences, the agent maintains a continuous and valid trajectory with the manipulated object, and the action generator smoothly adapts joint motions to the updated trajectories. These results highlight DecHOI’s ability to sustain collision-free and consistent coordination between human and object across extended dynamic environments.
4.6 Ablation Study
Condition Matching Human Motion Quality Interaction Quality
Methods $T_{s}\downarrow$ $T_{e}\downarrow$ $FS\downarrow$ $R_{prec}\uparrow$ $C_{F1}\uparrow$ $P_{\text{hand}}\downarrow$ $P_{\text{body}}\downarrow$
Baseline 1.72 7.92 0.35 0.67 0.65 0.58 0.60
w/ adversarial 1.68 7.78 0.39 0.64 0.66 0.53 0.57
w/ cross-attention 1.75 7.82 0.38 0.70 0.54 0.56 0.56
Ours (DecHOI) 1.59 6.91 0.38 0.72 0.67 0.53 0.54
Table 4: Ablation results for DecHOI on the FullBodyManipulation [li2023object], evaluating the contribution of each component.
As shown in Tab. 4, we perform an ablation study on the FullBodyManipulation [li2023object] to analyze the effects of adversarial training and text conditioning based on cross-attention. The baseline, which uses only the decoupled generative modeling and concatenates text embeddings along the sequence axis following prior work [li2024controllable, wu2025human], achieves moderate performance but suffers from high penetration errors and limited contact consistency. Introducing adversarial training substantially reduces $P_{\text{hand}}$ and $P_{\text{body}}$, indicating improved realism and more stable interactions, although the linguistic alignment measured by $R_{\text{prec}}$ slightly decreases. Conversely, adding cross-attention layers for text conditioning enhances $R_{\text{prec}}$ by effectively capturing semantic intent, but slightly degrades interaction quality due to weaker regularization. When both modules are combined in DecHOI, the model achieves the best overall performance, simultaneously improving $R_{\text{prec}}$, $C_{F1}$, and penetration metrics. These results indicate that adversarial training enhances the robustness of distal joint interactions during object manipulation, and that text conditioning based on cross-attention enables more explicit transfer of semantic information across modalities, leading to stronger alignment.
4.7 User Study
Figure 8: Stacked horizontal bars showing user preference distributions for DecHOI compared with CHOIS [li2024controllable] and HOIFHLI [wu2025human] across two evaluation criteria: Text Alignment and Interaction Quality.
We conducted a user study to evaluate the perceptual realism and text alignment of DecHOI compared with CHOIS [li2024controllable] and HOIFHLI [wu2025human]. For each comparison, a total of 60 text-scene pairs, 30 per method, were randomly sampled from our full test set, and 200 participants were recruited via Amazon Mechanical Turk (AMT). Each participant viewed two anonymized clips, each lasting about four seconds, in random order and answered two questions: i) which video better matches the text instruction (Text Alignment), and ii) which video shows better human-object interaction quality (Interaction Quality). Responses were collected using a two-choice scale: Prefer Ours or Prefer Others. As summarized in Fig. 8, DecHOI was consistently preferred on both criteria across both baselines, confirming that our decoupled framework yields motions that are semantically faithful to textual instructions and more realistic, stable, and natural than prior methods.
5 Conclusion
In this work, we propose a decoupled generative modeling framework for human-object interaction synthesis that separates trajectory generation from fine-grained action generation. This formulation lowers optimization complexity and enables the model to learn precise and stable motion, which reduces temporal asynchrony and penetration. It also removes the need for manually specified intermediate waypoints, improving practical applicability. Additionally, a compact adversarial discriminator focused on distal joint cues further improves interaction fidelity and strengthens contact realism. We also present a dynamic planner that delivers robust and adaptive path planning in long-horizon, multi-agent settings, as demonstrated on the DynaPlan. Extensive quantitative and qualitative results show state-of-the-art performance across major metrics, producing realistic and semantically aligned human-object interaction.